# PlainQAFact: Retrieval-augmented Factual Consistency Evaluation Metric for Biomedical Plain Language Summarization

### v3

@@ Appendix A Dataset Annotation Protocol
We developed a comprehensive annotation procedure for freelancers on Upwork to conduct fact-checking annotations.
Our annotation procedure involves two stages, starting with thorough training using detailed examples to ensure consistent understanding of the task between annotators. Annotators receive a spreadsheet where each row contains a pair of data: a sentence extracted from the plain language summary and its corresponding scientific abstract. For every sentence-abstract pair, the annotators are required to label three features: External, Category, and Relation, with the appropriate labels.
@@
1. External: Determine whether the sentence includes information does not present in the scientific abstract. 1) Yes: The sentence contains external information that is not explicitly mentioned, paraphrased, or implied in the scientific abstract. 2) No: The sentence contains information that is explicitly stated or closely paraphrased from the scientific abstract. 2. Category: Classify the sentence of the plain language summary into one of the following categories (you can only choose one category per annotation) 1) Definition: Provides a fundamental explanation of a term. 2) Background: Information that helps understand the term within the context of the abstract, such as relevance, significance, or motivation. 3) Example: Specific instances that illustrate the use of the term in the scientific abstract. 4) Method/result: Details about the methodology or results described in the scientific abstract. 5) Other: For sentences that do not fit into the categories above, please indicate the category 3. Relation: Identify the sentence(s) in the scientific abstract that the sentence of the plain language summary is related to. Use indices of sentences from the scientific abstract to link the sentence of the plain language summary. You can select one or more sentences from the scientific abstract. List the indices like s1_1,s2_3,s3_6. If no relation is found, mark it as “external.”
@@
Each annotator will annotate 40 summary-abstract pairs to ensure each sentence of the plain language summary has two sets of annotations from different people. For each row of the spreadsheet, they need to annotate three columns: “External,” “Category,” and “Relation.”
Annotation Fields to Be Completed <<TO BE ANNOTATED>> External: Mark “Yes” or “No” to indicate if the sentence in the ‘Target_Sentence’ column contains external information.
<<TO BE ANNOTATED>> Category: Choose the most fitting category of the sentence from the list (Definition, Background, Example, Method/Result, Other).
<<TO BE ANNOTATED>> Relation: List the relevant sentence indices from the abstract in the ‘Original_Abstract’ column that relate to the plain language sentence (e.g., s10_1,s10_5). For example, filling in ‘s10_1,s10_5,s10_9’ if you think these three sentences from the abstract are relevant to sentence t1_3. Use commas to separate multiple indices.
@@ Appendix B Dataset Examples
Category Background
Relation external
@@
Category Background
Relation s10_1,s10_2
@@
Category Method/Result
Relation s15_16,s15_17,s15_20
@@
In Table 7, we presents three representative examples from PlainFact. Each example is annotated with five features: a plain language sentence, an indicator of whether the sentence is simplification or explanation, its category, its relation, and the corresponding abstract. All plain language sentences are factual, as they were written by the authors from the Cochrane database. “Need External Information?” feature specifies whether a sentence can be validated solely by the abstract. A “Yes” label indicates that the sentence includes information not explicitly mentioned in the abstract, and vice versa. The “Relation” feature identifies the sentence(s) in the scientific abstract most relevant to the plain language summaries; if no corresponding content exists in the abstract, the relation is marked as “external.” We randomly sample three examples from the dataset to to illustrate the dataset’s structure. Additionally, indexes have been created for both the plain language sentences and the abstract sentences to facilitate annotation.
@@ E.4 LLM Prompts
We utilize two types of prompts to guide LLMs through two stages of factual consistency evaluation: classification and answer extraction. In accordance with the benchmark annotation protocol (A), we employ GPT-4o as a classifier to determine whether a given sentence or summary requires elaborative explanations for factual verification. For both stages, we set the max_tokens parameter to 512 and configure the temperature to 0 for GPT-4o and 0.01 for the Llama 3.1 8B Instruct model.
@@
For the AE stage, we explore both GPT-4o and Llama 3.1 8B Instruct as backbone models to assess the factual consistency evaluation performance of PlainQAFact. Following the task description outlined in QAFactEval [39], we instruct both LLMs to extract potential answer entities from the input PLS.
@@
Example 3: Prompt of Llama 3.1 8B Instruct as the Answer Extractor.
Additionally, we report the performance of using Llama 3.1 8B Instruct as a judge to evaluate the factual consistency of a given PLS based on its source scientific abstract.
@@
Example 4: Prompt of Llama 3.1 8B Instruct as a factual consistency judge.

### v4

@@ Appendix A Dataset Annotation Protocol
We develop a comprehensive annotation procedure for freelancers on Upwork to conduct fact-checking annotations.
Our annotation procedure involves two stages, starting with thorough training using detailed examples to ensure consistent understanding of the task between annotators. Annotators receive a spreadsheet where each row contains a pair of data: a sentence extracted from the plain language summary and its corresponding scientific abstract. For every sentence-abstract pair, the annotators are required to label three features: Factuality type, Functional role, and Sentence alignment, with the appropriate labels.
@@
1. Factuality type: Determine whether the sentence includes information does not present in the scientific abstract. 1) Yes: The sentence contains external information that is not explicitly mentioned, paraphrased, or implied in the scientific abstract. 2) No: The sentence contains information that is explicitly stated or closely paraphrased from the scientific abstract. 2. Functional role: Classify the sentence of the plain language summary into one of the following categories (you can only choose one category per annotation) 1) Definition: Provides a fundamental explanation of a term. 2) Background: Information that helps understand the term within the context of the abstract, such as relevance, significance, or motivation. 3) Example: Specific instances that illustrate the use of the term in the scientific abstract. 4) Method/result: Details about the methodology or results described in the scientific abstract. 5) Other: For sentences that do not fit into the categories above, please indicate the category 3. Sentence alignment: Identify the sentence(s) in the scientific abstract that the sentence of the plain language summary is related to. Use indices of sentences from the scientific abstract to link the sentence of the plain language summary. You can select one or more sentences from the scientific abstract. List the indices like s1_1,s2_3,s3_6. If no relation is found, mark it as “external.”
@@
Each annotator will annotate 40 summary-abstract pairs to ensure each sentence of the plain language summary has two sets of annotations from different people. For each row of the spreadsheet, they need to annotate three columns: “Factuality type,” “Functional role,” and “Sentence alignment.”
Annotation Fields to Be Completed <<TO BE ANNOTATED>> Factuality type: Mark “Yes” or “No” to indicate if the sentence in the ‘Target_Sentence’ column contains external information.
<<TO BE ANNOTATED>> Functional role: Choose the most fitting category of the sentence from the list (Definition, Background, Example, Method/Result, Other).
<<TO BE ANNOTATED>> Sentence alignment: List the relevant sentence indices from the abstract in the ‘Original_Abstract’ column that relate to the plain language sentence (e.g., s10_1,s10_5). For example, filling in ‘s10_1,s10_5,s10_9’ if you think these three sentences from the abstract are relevant to sentence t1_3. Use commas to separate multiple indices.
@@ Appendix B Dataset Examples
Functional role Background
Sentence alignment external
@@
Functional role Background
Sentence alignment s10_1,s10_2
@@
Functional role Method/Result
Sentence alignment s15_16,s15_17,s15_20
@@
In Table 7, we presents three representative examples from PlainFact. Each example is annotated with five features: a plain language sentence, an indicator of whether the sentence is simplification or explanation, its category, its relation, and the corresponding abstract. All plain language sentences are factual, as they were written by the authors from the Cochrane database. “Need External Information?” feature specifies whether a sentence can be validated solely by the abstract. A “Yes” label indicates that the sentence includes information not explicitly mentioned in the abstract, and vice versa. The “Sentence alignment” feature identifies the sentence(s) in the scientific abstract most relevant to the plain language summaries; if no corresponding content exists in the abstract, the relation is marked as “external.” We randomly sample three examples from the dataset to to illustrate the dataset’s structure. Additionally, indexes have been created for both the plain language sentences and the abstract sentences to facilitate annotation.
@@ E.4 LLM Prompts
We utilize two types of prompts to guide LLMs through two stages of factual consistency evaluation: classification and answer extraction. In accordance with the benchmark annotation protocol (A), we employ GPT-4o as a classifier to determine whether a given sentence or summary requires elaborative explanations for factual verification. For both stages, we set the max_tokens parameter to 512 and configure the temperature to 0 for GPT-4o and 0.01 for the Llama 3.1 model.
@@
For the AE stage, we explore both GPT-4o and Llama 3.1 as backbone models to assess the factual consistency evaluation performance of PlainQAFact. Following the task description outlined in QAFactEval [39], we instruct both LLMs to extract potential answer entities from the input PLS.
@@
Example 3: Prompt of Llama 3.1 as the Answer Extractor.
Additionally, we report the performance of using Llama 3.1 as a judge to evaluate the factual consistency of a given PLS based on its source scientific abstract.
@@
Example 4: Prompt of Llama 3.1 as a factual consistency judge.